June 22, 2022 Calorimeter
Meeting Day/Time: every other WEDNESDAY at 11:00 a.m. JLab time
Use ZoomGov: https://jlab-org.zoomgov.com/j/1608552106?pwd=UE1nZFIzUUZHVjJEVGJXTGdUYTV3UT09.
Our meeting ID number is 160 855 2106. Click "Expand" for more details.
- Meeting ID: 160 855 2106 Passcode: 944454
- One tap mobile
+16692545252,,1608552106# US (San Jose) +16468287666,,1608552106# US (New York)
- Dial by your location
+1 669 254 5252 US (San Jose) +1 646 828 7666 US (New York) +1 551 285 1373 US +1 669 216 1590 US (San Jose) 833 568 8864 US Toll-free
- Meeting ID: 160 855 2106
- Find your local number: https://jlab-org.zoomgov.com/u/acAwPdiE89
- Join by SIP email@example.com
- Join by H.323
184.108.40.206 (US West) 220.127.116.11 (US East) Meeting ID: 160 855 2106 Passcode: 944454
- FCAL HDFCAL log book
- BCAL HDBCAL log book
- CCAL HDCCAL log book
- Calibrations: 2020 Data Production; RunPeriod-2019-11 Validation; Offline Monitoring Data Validation
- Photon Reconstruction Task Force
- BCAL calibration links, 2020, PrimEx, SRC
Goals for Calorimetry Group
- Determine preliminary photon reconstruction efficiencies as a function of E, phi and theta in data and simulation with a point-to-point precision of at least 5%.
- Measure systematics of pi0/eta mass calibration as a function of detector position to a precision of at least 5 MeV.
- Demonstrate agreement of photon reconstruction efficiency and resolution between data and simulation as a function of E, phi and theta to within 5%.
- Short term CALWG Items
- Long term items
- Calorimeter work packages
- AI - BCAL Monitoring:
- Past Meeting Minutes
- Run Operations
- BCAL efficiency: Churamani 
- Any other business
Attendees:Igal, Kevin Saldana, Varun, Stjepan, Jon, Karthik, Malte, Churamani, Tolga, Sean, Elton, Beni, Rebecca
- Welcome to Kevin (IU grad student)
- CPP run is going ok.
- Zisis is on vacation in July. Malte volunteered to assist Mark in organizing and running the CALWG meetings. Jon can also help.
- Past Meeting Minutes: no review.
- Run Operations
- FCAL: Malte provided an update on the base replacement. Failure rate is much lower than last year; preventative maintenance has likely paid off. The pool of spares is 30, with 15 ready to go. The 30 should last until the end of CPP and Chris is working on more boards (5-10/week) which then need a 2-week testing period. Also, there have been 0 lockups (!) so far n CPP. The RL is about the same in CPP as LH2 (5%) and a bit lower than PrimEx (8%). Rate in FCAL is lower during CPP and this might be helping with no lockups.
- BCAL: nominal operation. (See LED monitoring report below).
- CPP update: Igal presented an update on CPP; see FCAL, CPP FCAL time/energy calibrations, Gain matching calibration plots). General comments: pi0 is not visible for some runs; these have low statistics but there could be other issues. The eta is not visible at all in stark contrast to the SRC runs with 12C. Simon is working on the vertex which is showing strange behaviour. Timing calibrations (e.g. TOF) may all need to be refined. These adjustments may help.
- CCAL: no report.
- Sean revisited FCAL calibration of GlueX-I data. The issue is whether to use Igal's PrimEx method for the calibration of the FCAL inner blocks and on what timescale. A quick calibration could be done in 1-2 weeks on the 2017 data with the existing clutserizer. LOG vs COG weighting was discussed; it is easy to flip form one to the other. Implementation of a different algorithm (e.g. Island) would take about 1 year to implement.
- BCAL: Stjepan gave an update on the status of the SRC and PrimEx BCAL gain calibrations, including non-linear energy correction. Things look quite stable. It was decided to have the SRC constants uploaded to ccdb, which Stjepan and Karthik will do. Sean suggested that Mark be contacted to ask whether for PrimEx the low level constants were sufficiently different to warrant Stjepan to rerun the gain calibrations.
- BCAL LED: Varun gave a report on the LED monitoring. Upstream trigger bit is missing; past remedy was to reboot the crate. Mark does this. SiPMs look nominal.
- BCAL (offline): Tolga gave an update on the hadronic Efficiencies plots for Version 7 Monitoring. A discussion ensued on what is expected in Layer 4 (lower efficiency mostly because particles stop in Layer 3). Tolga will look at the run sheets and determine which runs are good, which should be kept even if a handful of channels are hot, and which should be rejected. Elton will assist in the interpretation of the efficiency results.
- FCAL: Kevin gave a presentation on the FCAL Timing Resolution, using the LED system. The overall timing resolution looks good. See link to slides above.
- BCAL efficiency: Churamani gave an update on the Tag and Probe method using the eta meson. Cuts are the same as before except for chi2/ndg which is now >10. A discussion took place on the fiducial cut in terms of polar angle: 11 deg is too close to where we can not reliably reconstruct photons in the BCAL. Work is ongoing.
- Any other business